OpenStudy (anonymous):

MEDAL 1. Antony works at a local clothing store. He can earn either $600 per month plus a 5% commission or $450 per month and a 7% commission. Write and solve a system of equations that you could use to determine the amount of merchandise that Antony would need to sell in order to earn the same amount of money for each pay scale.

OpenStudy (danjs):

600 + 0.05 *x = 450 + 0.07*x

OpenStudy (danjs):

where x is the amount of sales

OpenStudy (danjs):

Earnings = $600 + 0.05*(Sales) Or Earnings = $450 + 0.07*(Sales)

OpenStudy (danjs):

Sales = $7500 to earn the same amount in both cases

OpenStudy (danjs):

Earnings in Case 1 = $600 + 0.05*(Sales) Earnings in Case 2 = $450 + 0.07*(Sales) We want, Case 1 = Case 2 600 + 0.05 *Sales = 450 + 0.07*Sales Sales = 7500 dollars